OpenStudy (anonymous):

2x + 3y = 1200 Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (anonymous):

this is the graph.

OpenStudy (anonymous):

@jim_thompson5910 @Johnbc @jdoe0001 @THEBEAST2198

OpenStudy (anonymous):

Solve for Y, meaning isolate y by itself so that you may have F(x) or Y as a function of x.

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

is it where you break down the equation to where you get y=400?

OpenStudy (anonymous):

No it is where you just use algebra to get all the variables of Y on the left and the rest on the right or vice versa.

OpenStudy (jdoe0001):

seems to me that's a dead horse you might be beating up upon

OpenStudy (anonymous):

@jdoe0001 Has provided a very nice explanation of how you might go about graphing as well as describing how to graph. Please refer to your earlier post.

OpenStudy (anonymous):

can you just explain it on here, its not opening my old post. :(

OpenStudy (anonymous):

2x 3y = 1200. Subtract 2x from both sides: 2x -2x 3y = 1200 -2x Now Divide both sides by the number next to y. In this case 3: 3y/3 = 1200/3 -2x/3 Now we're left with y = 400 -2x/3 Move variables to the left y = -2x/3 400. @MonikaCristello

OpenStudy (anonymous):

oh wait nvm -_____-